OpenStudy (anonymous):

Please Help will medal!! −e^(4−5k) +5.2=−6

OpenStudy (anonymous):

\[\large −e^{4−5k} +5.2=−6\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

subtract \(5.2\) change the sign write in log form solve for \(k\)

OpenStudy (anonymous):

ie start with \[4-5k=\ln(11.2)\]

OpenStudy (anonymous):

You have to do the same trick as the previous question I answered. Get whatever you can to the other side with -6. Keep e^(4-5k) Treat the exponent of e as the end result of the ln of the value you got at the other side of the = sign. Then use that solution to set up the equation 4-5k = the result of ln (number) Evaluate k